# ROGR: Relightable 3D Objects using Generative Relighting

### v1

@@ 5.1 TensoIR benchmark
  NVDiffRecMC [57] $\dagger$   25.08   32.28   0.974   0.027
  NVDiffRec [15] $\dagger$   24.93   32.42   0.975   0.027
  PhySG [54]   21.81   28.11   0.960   0.055
  NVDiffRec [15]   22.91   29.72   0.963   0.039
  NeRD [19]   23.29   29.65   0.957   0.059
  NeRFactor [52]   23.54   30.38   0.969   0.048
  InvRender [53]   23.76   30.83   0.970   0.046
  NVDiffRecMC [57]   24.43   31.60   0.972   0.036
  Neural-PBIR [55]   26.01   33.26   0.979   0.023
  R3DG [56]   21.25   27.50   0.962   0.063
  Neural Gaffer [6]   23.16   29.94   0.966   0.047
  IllumiNeRF [7]   25.56   32.74   0.976   0.027
@@
Table 2: StandfordORB benchmark [48]. We evaluate fourteen objects captured in the real world. Each object was captured in three different lightings. For each object-lighting pair, we evaluate novel view renderings under the other two lighting. The benchmark contains 840 renderings in total. $\dagger$ denotes models trained with the ground-truth 3D scans and pseudo materials optimized from light-box captures. Best and 2nd-best are highlighted.
@@
Figure 5: Qualitative comparisons on Stanford-ORB [48]. Renderings from all methods are rescaled to the image resolution of the ground truth. Compared to previous work, our method produces high-fidelity renderings with more faithful specular reflections highlighted in the red boxes.
@@ 5.2 Stanford-ORB benchmark
In Fig. 5, our method also achieves state-of-the-art performance on the Stanford-ORB dataset, and is most effective at reflective objects like the “baking” and “ball” scenes where consistent reflections clearly show up in the reconstruction. We provide quantitative comparisons in Tab. 2, where our method outperforms others in PSNR-H and SSIM, and acheive second-best results in PSNR-L and LPIPS. As discussed in IllumiNeRF [7], our results are also qualitatively superior to those of Neural-PBIR [55], but they are worse quantitatively due to the mostly-diffuse renderings of Neural-PBIR.
@@ 5.4 Ablation Studies
  Ours full model, 4 view, 111 envmaps   31.04   0.86   0.082
  Ours full model, 16 view, 111 envmaps   31.86   0.90   0.077
  Our full model, 64-view, 10 envmaps   29.12   0.82   0.086
  Our full model, 64-view, 50 envmaps   31.78   0.88   0.079
  Our full model, 64-view, 150 envmaps   31.90   0.92   0.075
Table 3: Ablation studies on the “hotdog” scene from TensoIR [45]. See the text and Fig. 6 for corresponding qualitative results and additional explanations. Best and 2nd-best are highlighted.
@@
Our multi-scale specular conditioning features, which are computed by blurring the environment map according to Eq. 4 using multiple kernel sizes, are provided to the model. Each scale is designed to approximate the incident light averaged over a set of directions corresponding to a particular surface roughness. When we skip this blurring operation and use the environment map values directly, our model can still represent highly specular regions (see Fig. 6(a)), but struggles more with rough or diffuse surfaces, leading to a drop in reconstruction metrics.
@@
An alternative to our conditioning signals is a per-image appearance embedding vector, similar to the GLO codes in NeRF-in-the-Wild [42]. While this allows the model to be trained on multiple illuminations, it does not generalize to new unseen lights. Additionally, unlike our model, we found that the embedding vectors do not scale well to a large number of illumination conditions, resulting in significantly worse qualitative and quantitative results.
@@
We study the effect of environment map count on relightable NeRF training by using 10, 50, 111, and 150 illuminations. As shown in Tab. 3, more illuminations generally improve relighting performance, which saturates around 111 maps.
@@ Appendix A Supplementary Webpage and Videos
We suggest readers check the websiter in our supplementary material. We provide more video renderings, including ablation studies and result comparisons on TensoIR [22], StanfordORB [48], and real-world objects.
@@ B.1 Multi-view Relighting Diffusion
We implement our multi-view relighting diffusion model using JAX [58]. It is initialized from a pre-trained latent diffusion model for text-to-image generation, similar to StableDiffusion [51]. Our model denoises multiple noisy latents of size $64\times 64\times 8$ and decodes them into multi-view relit images of size $512\times 512\times 3$. Since our model is not conditioned on text prompt, we only feed empty strings to the CLIP text encoder [59].
@@
In Fig. 7, we provide the detailed inputs of our relighting diffusion model. While our diffusion model receives and produces image latents, we depict them as images for better clarity. We encode RGB images under source lighting, HDR, and LDR target environment maps into latents of size $64\times 64\times 8$. The raymaps consist of ray origins and ray directions corresponding to image pixels. We downscale them from the original image resolution $512\times 512\times 6$ to the latent size $64\times 64\times 6$. We concatenate the source image latent, HDR, LDR environment latents, raymaps, and noisy latents along the channel dimension to form a new feature of size $64\times 64\times 38$, and then feed it into a multi-view diffusion network to produce clean target latents of size $64\times 64\times 8$. Our denoiser network is based on CAT3D [37]. Please refer to Fig. 7 of CAT3D for the network architecture details.
During training, we use the DDPM schedule, with beta values that linearly increase from $8.5\times 10^{-4}$ to $1.2\times 10^{-2}$ over 1024 steps. We use noise prediction as our diffusion objective. The model was trained on 128 TPU v5 chips using a learning rate of $10^{-4}$, with a total batch size of 128 for 360k iterations and 10K warm-up steps. We adopted a progressive training scheme, where we first trained a 4-view diffusion model for 300k steps, and then fine-tuned it for 16-view diffusion for 15k steps, and finally fine-tuned it for 64-view diffusion for 45 steps. We keep the learning rate as $10^{-4}$ when we fine-tune the model to relight the large number of views. We enable classifier-free guidance (CFG) [60] by randomly dropping the HDR and LDR environment maps with a probability of 0.1. During inference, we use the DDIM schedule [61] with 50 sampling steps and the classifier-free weight is set to 3.0.
@@ B.2 Relightable NeRF
Our relightable NeRF is mainly based on NeRF-Casting [40], which implicitly learnt the accumulated reflection features to model accurate and detailed reflections. Instead, we choose to explicitly learn these from given enviroment maps. Given a single ray with origin $\mathbf{o}$ and direction $\mathbf{d}$, we sample $N$ points $\mathbf{x}^{(i)}$ along the ray and use multi-resolution hash grid and MLP to encode $\mathbf{x}^{(i)}$ into density $\mathbf{\tau}^{(i)}$, roughness $\mathbf{\rho}^{(i)}$, and surface normal $\mathbf{n}^{(i)}$. Then, they are alpha composited to compute a single expected termination point $\mathbf{\overline{x}}$, a von Mishes-Fisher distribution (vMF) width $\overline{k}$, and surface normal $\mathbf{\overline{n}}$. Next, we construct a reflection cone by reflecting $\mathbf{d}$ around the micro-surface to obtain a vMF distribution over reflected rays $vMF(\mathbf{d}^{\prime},\overline{k})$. We sample $K=5$ reflected rays with location $\mathbf{o}^{\prime}$ and $\mathbf{d}^{\prime}_{j}$. We cast these rays and sample $N^{\prime}$ points $\mathbf{x}^{(i)}_{j}$ along each reflection ray. $N^{\prime}$ points $\mathbf{x}^{(i)}_{j}$ are encoded into $N^{\prime}$ densities. Based on location $\mathbf{o}^{\prime}$ and $\mathbf{d}^{\prime}_{j}$, we can query illumination information from environment maps and encode them with $\mathbf{x}^{(i)}_{j}$ into features $\mathbf{f}^{(i)}_{j}$ through an MLP. These features $\mathbf{f}^{(i)}_{j}$ are alpha-composed along each ray to get per-ray reflection feature ${\mathbf{\overline{f}}^{(i)}}_{j}$, which can be further averaged into a single reflection feature $\mathbf{f}$. Finally, $\mathbf{f}$, bottleneck geometry feature $\mathbf{b}^{(i)}$, mixing coefficients $\beta^{(i)}$ and view direction $\mathbf{d}$ are feed into color decoder to predict the color value of $\mathbf{x}^{(i)}$.
@@ C.1 Training data preprocessing
We use Objaverse [46] and an internal dataset containing high-quality 3D assets as our training data. The Objaverse dataset is released under the Apache-2.0 license.
To ensure high-quality renderings for training, we filter out low-quality 3D assets from Objaverse using the object list provided in [62], and further exclude (semi-)transparent objects, as our focus is on reflective and shiny surfaces. For 3D assets lacking texture or material information, we assign a uniform color sampled from $[0,1]$ as texture. We randomly sample three values from $[0,1]$ as the diffuse, roughness, and metallic terms of the material model.
Our relighting diffusion model requires multi-view images under diverse lighting conditions. To this end, we use 590 equirectangular environment maps from [47]. For each object, we randomly sample 64 camera poses on a sphere centered around the object. The camera distance ranges from $[0.5,2.0]$. For each camera view, we randomly select 16 environment maps, augmenting each with a random horizontal shift. Then we use Blender’s Cycles path tracer to render images at a resolution of $\times512$, with 512 samples per pixel.
@@
We use TensoIR [22] and StanfordORB [48] datasets for relighting evaluations. TensoIR is under the MIT license. StanfordORB does not specify a license in the official GitHub repository.
TensoIR is a synthetic dataset that contains renderings of four objects under six lighting conditions. We use the train split of 100 views with “sunset” lighting condition as inputs for relightable NeRF. We then evaluate 200 novel views under other five environment maps, including “bridge”, “fireplace”, “forest”, “city”, and “night”. In total, we have 4,000 renderings for evaluation metric calculation.
Stanford-ORB is real-world benchmarks by data capture in the wild. It has 14 objects composed of various materials. Each object is captured under three distinct lighting conditions, producing a total of 42 (object, lighting) combinations. Following its evaluation protocol, we use images of an object under a single lighting condition and evaluate novel views under the two target lighting settings.
@@ C.3 Real-world data capture
We capture real-world objects to evaluate our relighting model using a Sony DSLR camera. Each object is placed on a table, and a handheld camera is used to record an image sequence by moving around the object. Each sequence contains approximately 160 to 300 frames. Camera intrinsics and extrinsics are estimated using COLMAP[63] for NeRF training. For foreground segmentation, we apply the pre-trained SAM [64] model. We will release our captured dataset upon publication under a non-commercial academic license.
@@ C.4 Evaluation metrics
We evaluate the relighting rendering quality by PSNR, SSIM [49], and LPIPS-VGG [50] for low dynamic range (LDR) images. On the Stanford-ORB benchmark, we also compute PSNR for high dynamic range (HDR) images, denoted as PSNR-H, while PSNR for LDR images is referred to as PSNR-L. For methods that only produce LDR images, we apply the inverse process of the sRGB tone mapping to transform the outputs into linear values. To address ambiguities in the relighting task, we align the predicted outputs with the ground truth images by applying channel-wise scale factors prior to metric computation. For Stanford-ORB, these scale factors are determined separately for each output image. For TensoIR, a single global scale factor is calculated and uniformly applied to all output images.
Appendix D Additional Results
D.1 Normal map visualization.
Appendix E Social Impacts

### v3

@@ 5.1 TensoIR benchmark
  NVDiffRecMC [NVDiffRecMC] $\dagger$   25.08   32.28   0.974   0.027
  NVDiffRec [NVDiffRec] $\dagger$   24.93   32.42   0.975   0.027
  PhySG [zhang2021physg]   21.81   28.11   0.960   0.055
  NVDiffRec [NVDiffRec]   22.91   29.72   0.963   0.039
  NeRD [boss2021nerd]   23.29   29.65   0.957   0.059
  NeRFactor [zhang2021nerfactor]   23.54   30.38   0.969   0.048
  InvRender [zhang2022modeling]   23.76   30.83   0.970   0.046
  NVDiffRecMC [NVDiffRecMC]   24.43   31.60   0.972   0.036
  Neural-PBIR [Neural-PBIR]   26.01   33.26   0.979   0.023
  R3DG [R3DG]   21.25   27.50   0.962   0.063
  Neural Gaffer [jin2024neural_gaffer]   23.16   29.94   0.966   0.047
  IllumiNeRF [zhao2024illuminerf]   25.56   32.74   0.976   0.027
@@
Table 2: StandfordORB benchmark [kuang2024stanford]. We evaluate fourteen objects captured in the real world. Each object was captured in three different lightings. For each object-lighting pair, we evaluate novel view renderings under the other two lighting. The benchmark contains 840 renderings in total. $\dagger$ denotes models trained with the ground-truth 3D scans and pseudo materials optimized from light-box captures. Best and 2nd-best are highlighted.
@@
Figure 5: Qualitative comparisons on Stanford-ORB [kuang2024stanford]. Renderings from all methods are rescaled to the image resolution of the ground truth. Compared to previous work, our method produces high-fidelity renderings with more faithful specular reflections highlighted in the red boxes.
@@ 5.2 Stanford-ORB benchmark
In Fig. 5, our method also achieves state-of-the-art performance on the Stanford-ORB dataset, and is most effective at reflective objects like the “baking” and “ball” scenes where consistent reflections clearly show up in the reconstruction. We provide quantitative comparisons in Tab. 2, where our method outperforms others in PSNR-H and SSIM, and acheive second-best results in PSNR-L and LPIPS. As discussed in IllumiNeRF [zhao2024illuminerf], our results are also qualitatively superior to those of Neural-PBIR [Neural-PBIR], but they are worse quantitatively due to the mostly-diffuse renderings of Neural-PBIR.
@@ 5.4 Ablation Studies
Table 3: Ablation studies on the “hotdog” scene from TensoIR [jin2023tensoir]. See the text and Fig. 6 for corresponding qualitative results and additional explanations. Best and 2nd-best are highlighted.
@@
Our multi-scale specular conditioning features, which are computed by blurring the environment map using multiple kernel sizes as in Eq. 4, are provided to the model. Each scale is designed to approximate the incident light averaged over a set of directions corresponding to a particular surface roughness. When we skip this blurring operation and use the environment map values directly, our model can still represent highly specular regions (see Fig. 6(a)), but struggles more with rough or diffuse surfaces, leading to a drop in reconstruction metrics.
@@
An alternative to our conditioning signals is a per-image appearance embedding vector, similar to the GLO codes in NeRF-in-the-Wild [martinbrualla2020nerfw]. While this allows the model to be trained on multiple illuminations, it does not generalize to new unseen lights. Additionally, unlike our model, we found that the embedding vectors do not scale well to a large number of illumination conditions, resulting in significantly worse qualitative and quantitative results.
@@
  Method   PSNR $\uparrow$   SSIM $\uparrow$   LPIPS $\downarrow$
  Ours full model, 4 view, 111 envmaps   31.04   0.86   0.082
  Ours full model, 16 view, 111 envmaps   31.86   0.90   0.077
  Our full model, 64-view, 10 envmaps   29.12   0.82   0.086
  Our full model, 64-view, 50 envmaps   31.78   0.88   0.079
  Our full model, 64-view, 111 envmaps   31.88   0.91   0.075
  Our full model, 64-view, 150 envmaps   31.90   0.92   0.075
Table 4: Ablation studies of number of views and illulimantions on the “hotdog” scene from TensoIR [jin2023tensoir].Best and 2nd-best are highlighted.
@@
We also study the effect of environment map count on relightable NeRF training by using 10, 50, 111, and 150 illuminations. As shown in Tab. 4, more illuminations generally improve relighting performance, which saturates around 111 maps.
Normal map visualization.
@@ Acknowledgement.
Appendix – ROGR: Relightable 3D Objects using Generative Relighting
@@ Appendix A Supplementary Webpage and Videos
We suggest readers check the website in our supplementary material. We provide more video renderings, including ablation studies and result comparisons on TensoIR [tensoir], StanfordORB [kuang2024stanford], and real-world objects.
@@ B.1 Multi-view Relighting Diffusion
We implement our multi-view relighting diffusion model using JAX [bradbury2018jax]. It is initialized from a pre-trained latent diffusion model for text-to-image generation, similar to StableDiffusion [rombach2022high]. Our model denoises multiple noisy latents of size $64\times 64\times 8$ and decodes them into multi-view relit images of size $512\times 512\times 3$. Since our model is not conditioned on text prompt, we only feed empty strings to the CLIP text encoder [radford2021learning].
@@
In Fig. 8, we provide the detailed inputs of our relighting diffusion model. While our diffusion model receives and produces image latents, we depict them as images for better clarity. We encode RGB images under source lighting, HDR, and LDR target environment maps into latents of size $64\times 64\times 8$. The raymaps consist of ray origins and ray directions corresponding to image pixels. We downscale them from the original image resolution $512\times 512\times 6$ to the latent size $64\times 64\times 6$. We concatenate the source image latent, HDR, LDR environment latents, raymaps, and noisy latents along the channel dimension to form a new feature of size $64\times 64\times 38$, and then feed it into a multi-view diffusion network to produce clean target latents of size $64\times 64\times 8$. Our denoiser network is based on CAT3D [gao2024cat3d]. Please refer to Fig. 7 of CAT3D for the network architecture details.
During training, we use the DDPM schedule, with beta values that linearly increase from $8.5\times 10^{-4}$ to $1.2\times 10^{-2}$ over 1024 steps. We use noise prediction as our diffusion objective. The model was trained on 128 TPU v5 chips using a learning rate of $10^{-4}$, with a total batch size of 128 for 360k iterations and 10K warm-up steps. We adopted a progressive training scheme, where we first trained a 4-view diffusion model for 300k steps, and then fine-tuned it for 16-view diffusion for 15k steps, and finally fine-tuned it for 64-view diffusion for 45 steps. We keep the learning rate as $10^{-4}$ when we fine-tune the model to relight the large number of views. We enable classifier-free guidance (CFG) [ho2022classifier] by randomly dropping the HDR and LDR environment maps with a probability of 0.1. During inference, we use the DDIM schedule [song2020denoising] with 50 sampling steps and the classifier-free weight is set to 3.0.
@@ B.2 Relightable NeRF
Our relightable NeRF is mainly based on NeRF-Casting [verbin2024nerf], which implicitly learnt the accumulated reflection features to model accurate and detailed reflections. Instead, we choose to explicitly learn these from given enviroment maps. Given a single ray with origin $\mathbf{o}$ and direction $\mathbf{d}$, we sample $N$ points $\mathbf{x}^{(i)}$ along the ray and use multi-resolution hash grid and MLP to encode $\mathbf{x}^{(i)}$ into density $\mathbf{\tau}^{(i)}$, roughness $\mathbf{\rho}^{(i)}$, and surface normal $\mathbf{n}^{(i)}$. Then, they are alpha composited to compute a single expected termination point $\mathbf{\overline{x}}$, a von Mishes-Fisher distribution (vMF) width $\overline{k}$, and surface normal $\mathbf{\overline{n}}$. Next, we construct a reflection cone by reflecting $\mathbf{d}$ around the micro-surface to obtain a vMF distribution over reflected rays $vMF(\mathbf{d}^{\prime},\overline{k})$. We sample $K=5$ reflected rays with location $\mathbf{o}^{\prime}$ and $\mathbf{d}^{\prime}_{j}$. We cast these rays and sample $N^{\prime}$ points $\mathbf{x}^{(i)}_{j}$ along each reflection ray. $N^{\prime}$ points $\mathbf{x}^{(i)}_{j}$ are encoded into $N^{\prime}$ densities. Based on location $\mathbf{o}^{\prime}$ and $\mathbf{d}^{\prime}_{j}$, we can query illumination information from environment maps and encode them with $\mathbf{x}^{(i)}_{j}$ into features $\mathbf{f}^{(i)}_{j}$ through an MLP. These features $\mathbf{f}^{(i)}_{j}$ are alpha-composed along each ray to get per-ray reflection feature ${\mathbf{\overline{f}}^{(i)}}_{j}$, which can be further averaged into a single reflection feature $\mathbf{f}$. Finally, $\mathbf{f}$, bottleneck geometry feature $\mathbf{b}^{(i)}$, mixing coefficients $\beta^{(i)}$ and view direction $\mathbf{d}$ are feed into color decoder to predict the color value of $\mathbf{x}^{(i)}$.
@@ C.1 Training data preprocessing
We use Objaverse [deitke2023objaverse] and an internal dataset containing high-quality 3D assets as our training data. The Objaverse dataset is released under the Apache-2.0 license.
To ensure high-quality renderings for training, we filter out low-quality 3D assets from Objaverse using the object list provided in [tang2024lgm], and further exclude (semi-)transparent objects, as our focus is on reflective and shiny surfaces. For 3D assets lacking texture or material information, we assign a uniform color sampled from $[0,1]$ as texture. We randomly sample three values from $[0,1]$ as the diffuse, roughness, and metallic terms of the material model.
Our relighting diffusion model requires multi-view images under diverse lighting conditions. To this end, we use 590 equirectangular environment maps from [zaal2021polyhaven]. For each object, we randomly sample 64 camera poses on a sphere centered around the object. The camera distance ranges from $[0.5,2.0]$. For each camera view, we randomly select 16 environment maps, augmenting each with a random horizontal shift. Then we use Blender’s Cycles path tracer to render images at a resolution of $\times512$, with 512 samples per pixel.
@@
We use TensoIR [tensoir] and StanfordORB [kuang2024stanford] datasets for relighting evaluations. TensoIR is under the MIT license. StanfordORB does not specify a license in the official GitHub repository.
@@ C.3 Real-world data capture
We capture real-world objects to evaluate our relighting model using a Sony DSLR camera. Each object is placed on a table, and a handheld camera is used to record an image sequence by moving around the object. Each sequence contains approximately 160 to 300 frames. Camera intrinsics and extrinsics are estimated using COLMAP[schonberger2016structure] for NeRF training. For foreground segmentation, we apply the pre-trained SAM [kirillov2023segment] model. We will release our captured dataset upon publication under a non-commercial academic license.
@@ C.4 Evaluation metrics
We evaluate the relighting rendering quality by PSNR, SSIM [SSIM], and LPIPS-VGG [LPIPS] for low dynamic range (LDR) images. On the Stanford-ORB benchmark, we also compute PSNR for high dynamic range (HDR) images, denoted as PSNR-H, while PSNR for LDR images is referred to as PSNR-L. For methods that only produce LDR images, we apply the inverse process of the sRGB tone mapping to transform the outputs into linear values. To address ambiguities in the relighting task, we align the predicted outputs with the ground truth images by applying channel-wise scale factors prior to metric computation. For Stanford-ORB, these scale factors are determined separately for each output image. For TensoIR, a single global scale factor is calculated and uniformly applied to all output images.
Appendix D Social Impacts